OpenStudy (anonymous):

Find the exact value by using a half-angle identity. sin five pi divided by twelve

OpenStudy (anonymous):

\[\sin (5\pi/12)\]

OpenStudy (anonymous):

@sleepyjess

OpenStudy (anonymous):

do you know if they will be online soon?

OpenStudy (asnaseer):

HINT:\[\frac{5\pi}{12}=\frac{6\pi}{12}-\frac{\pi}{12}\]

OpenStudy (anonymous):

i'm really bad at these so i dont know what to do with that

OpenStudy (anonymous):

@asnaseer can you help me more

OpenStudy (asnaseer):

what rules are you aware of with sin?

OpenStudy (anonymous):

the half angle identity is \[\sin(x/2) = +or- \sqrt{\frac{ 1-cosx }{ 2 }}\]

OpenStudy (anonymous):

@asnaseer

OpenStudy (asnaseer):

You will need to make use of these two identities here:\[\sin(\frac{\pi}{2}-\theta)=\cos(\theta)\]and:\[\cos(2\theta)=2\cos^2(\theta)-1\]you will also need to know some common sin/cos values, e.g.:\[\cos(30^{\circ})=\cos(\frac{\pi}{6})=\frac{\sqrt{3}}{2}\]

OpenStudy (anonymous):

so its